A 12- quart cooling system is checked and found to be filled with a solution of 40% antifreeze. The desired strength of the solution is 60% antifreeze. How many quarts of solution need to be drained and replaced with pure antifreeze to reach the desired strength?
Answer by jorel1380(3719)

12x.6=7.2
m+n=12
.40m+n=7.2
4m+10n=72
4m+4n=48
6n=24
n=4
You need to replace 4 qts. of 40% solution with pure anti-freeze for the mix.
8x.4=3.2
3.2+4=7.2
.6x12=7.2..